# Exploring the Effects of AI Nonverbal Emotional Cues on Human Decision Certainty in Moral Dilemmas

### v1

@@ 3.2. Materials
To align with the behavior pattern of the AI assistant, our criteria for selecting stories required clearly defined perpetrators and victims, allowing the AI assistant to express anger toward perpetrators and sadness toward victims reasonably. Based on this, we chose two scenarios. The first involves deciding between paying a ransom to a guerrilla group, which enables further violence, or refusing and resulting in a hostage’s death. The second involves choosing between betraying a colleague, leading to their execution by militants, or not betraying them, which results in the death of a busload of strangers. Each story has two versions: one where a proscriptive norm prohibits action with benefits outweighing the costs, and the other prescribes action with benefits smaller than the costs. The scenarios are shown in Table Exploring the Effects of AI Nonverbal Emotional Cues on Human Decision Certainty in Moral Dilemmas. The four scenarios appeared randomly for each participant during the task.
@@ 4. RESULTS
We include multiple factors to answer RQs. We begin by outlining the changes in certainty associated with decisions in dilemma scenarios, followed by analyzing data that reflect participants’ perceived AI and its influence on participants. Our results reveal that AI assistants with nonverbal emotional cues could influence participants’ decision certainty in moral dilemmas, especially when AI assistants take anger cues, which could answer RQ1. Moreover, AI assistants with nonverbal emotional cues could also affect participants’ perceptions of AI’s influence on human moral thinking and have apparent gender differences in the type of nonverbal cue, which could answer RQ2.
@@ 4.3.1. AI Influence on Human Moral Thinking
Two-way ANCOVA’s results shown in Figure 5 indicate that AI with AC or SC impacts participants’ moral thinking and has gender differences. We selected participants’ emotional state, utilitarianism, and age as control variables, with gender and nonverbal emotional cues as independent variables. The main effect of gender was not significant ($F(2,128)=0.069$, $p=0.933$), the main effect of nonverbal emotional cues was not significant ($F(2,128)=0.700$, $p=0.498$), and the interaction between gender and nonverbal emotional cues was significant ($F(3,128)=5.522$, $p<0.01$).
Since the number of subjects in other genders was too small ($n=3$) to allow statistical analysis, we took “Self-identified Male” and “Self-identified Female” for further analysis.
@@
Figure 5. Impact of nonverbal emotional cues on AI assistants’ influence on participants’ moral thinking. Significant differences are indicated between groups(*: p ¡ 0.05, **: p ¡ 0.01). Error bars represent 95% confidence intervals.
@@
We find a significant effect of AC on self-identified males perceived AI assistants influence Moral Thinking than SC ($M_{AC}=2.024$, $SD=0.280$, $M_{SC}=0.592$, $SD=0.332$, $p<0.01$), but there was no significant difference between AC and the baseline ($M_{Baseline}=1.374$, $SD=0.261$, $p=0.254$) on that. Besides, there’s a significant effect of SC on Self-identified females perceived AI assistants influence Moral Thinking than the baseline ($M_{SC}=1.895$, $SD=0.265$, $M_{Baseline}=0.850$, $SD=0.271$, $p<0.05$), but there was no significant difference between SC and AC ($M_{AC}=0.990$, $SD=0.312$, $p=0.086$) on that.
The results show that self-identified male participants may perceive AI assistants taking AC rather than SC as nonverbal emotional cues having tremendously apparent effects on moral thinking. Besides, self-identified female participants may perceive AI assistants taking SC rather than the baseline as exceedingly noticeable in influencing moral thinking.
@@
We also find that in AC, the perception of AI assistants’ influence on moral thinking in self-identified males is significantly higher than that in self-identified females ($p<0.05$). However, in SC, the perception reflected in self-identified males is significantly lower than in self-identified females ($p<0.01$).
@@ 5.2. The Interaction Effect of Gender and Emotion on Perception
Our results shows that self-identified female participants may perceive AI assistants taking SC rather than baseline, while self-identified male participants may perceive AI assistants taking AC rather than SC. However, regardless of how they perceive the AI assistant, the reversal effect on certainty by the AI assistant is the same for participants of both genders, with the reversal effect of AC being more significant. This is an interesting phenomenon where people’s perception of AI does not align with the actual impact it has on them. The reasons behind this discrepancy in perception are worth exploring.
There is a stereotype that different genders tend to express different emotions. Studies have found that emotions such as sadness and depression are often seen as associated with women, while anger and aggression are generally considered to be more closely associated with men (Simon and Nath, 2004; Barrett and Bliss-Moreau, 2009; Craig and Lee, 2020; Plante et al., 2000). Building on our research, which found that participants of different genders perceive AI emotions differently, it raises the question of whether this phenomenon is related to the aforementioned stereotype. In other words, do the emotions that men and women are stereotypically expected to express also influence how they perceive emotions? This is an area that can be further explored.
@@ 5.4. Limitations and Future Opportunities
Second, our study focused exclusively on AI assistants designed to challenge users’ decisions without including AI that supports users or provides detailed analyses. This limitation prevents us from assessing how different AI roles influence user behavior and decision-making. Future research could compare challenging and supportive AI across various scenarios to identify the most appropriate approach for each context.
@@ 6. CONCLUSION
This study empirically examines the influence of nonverbal emotional cues from AI assistants on decision certainty and how humans perceive the influence of AI in moral dilemmas. Our findings demonstrate that agents expressing anger significantly induce reversal shifts in decision certainty. The interaction effects between participant gender and agents’ nonverbal emotional cues notably affect participants’ perceptions of AI’s influence. These results highlight the potential of subtle nonverbal emotional cues to shape human moral decisions, offering insights for designing future human-AI systems. We hope our findings contribute to ongoing discussions in the AI ethics community about the nuanced ways AI can impact decision-making processes.

### v2

# Anger Speaks Louder? Exploring the Effects of AI Nonverbal Emotional Cues on Human Decision Certainty in Moral Dilemmas

@@ 3.2. Materials
To align with the behavior pattern of the AI assistant, our criteria for selecting stories required clearly defined perpetrators and victims, allowing the AI assistant to express anger toward perpetrators and sadness toward victims reasonably. Based on this, we chose two scenarios. The first involves deciding between paying a ransom to a guerrilla group, which enables further violence, or refusing and resulting in a hostage’s death. The second involves choosing between betraying a colleague, leading to their execution by militants, or not betraying them, which results in the death of a busload of strangers. Each story has two versions: one where a proscriptive norm prohibits action with benefits outweighing the costs, and the other prescribes action with benefits smaller than the costs. The scenarios are shown in Table Anger Speaks Louder? Exploring the Effects of AI Nonverbal Emotional Cues on Human Decision Certainty in Moral Dilemmas. The four scenarios appeared randomly for each participant during the task.
@@ 4. RESULTS
We include multiple factors to answer RQs. We begin by outlining the changes in certainty associated with decisions in dilemma scenarios, followed by analyzing data that reflect participants’ perceived AI and its influence on participants. Our results reveal that the AI assistant with nonverbal emotional cues could influence participants’ decision certainty in moral dilemmas, especially when the AI assistant take anger cues, which could answer RQ1. Moreover, the AI assistant with nonverbal emotional cues could also affect participants’ perceptions of AI’s influence on human moral thinking and have apparent gender differences in the type of nonverbal cue, which could answer RQ2.
@@ 4.3.1. AI Influence on Human Moral Thinking
Two-way ANCOVA’s results shown in Figure 5 indicate that AI with AC or SC impacts participants’ moral thinking and has gender differences. We selected participants’ emotional state, utilitarianism, and age as control variables, with gender and nonverbal emotional cues as independent variables. The main effect of gender was not significant ($F(2,128)=0.069$, $p=0.933$), the main effect of nonverbal emotional cues was not significant ($F(2,128)=0.700$, $p=0.498$), and the interaction between gender and nonverbal emotional cues was significant ($F(3,128)=5.522$, $p<0.01$). Since the number of subjects in other genders was too small ($n=3$) to allow statistical analysis, we took “Self-identified Male” and “Self-identified Female” for further analysis.
@@
Figure 5. Impact of nonverbal emotional cues on the AI assistant’s influence on participants’ moral thinking. Significant differences are indicated between groups(*: p ¡ 0.05, **: p ¡ 0.01). Error bars represent 95% confidence intervals.
@@
We found a significant effect of AC on the perception of the AI assistant’s influence on moral thinking of self-identified males, compared to SC.($M_{AC}=2.024$, $SD=0.280$, $M_{SC}=0.592$, $SD=0.332$, $p<0.01$), but there was no significant difference between AC and the baseline ($M_{Baseline}=1.374$, $SD=0.261$, $p=0.254$) on that. Besides, there’s a significant effect of SC on Self-identified females’ perception of the AI assistant’s influence on moral thinking than the baseline ($M_{SC}=1.895$, $SD=0.265$, $M_{Baseline}=0.850$, $SD=0.271$, $p<0.05$), but there was no significant difference between SC and AC ($M_{AC}=0.990$, $SD=0.312$, $p=0.086$) on that.
The results show that self-identified male participants may perceive the AI assistant’s influence, when displaying AC rather than SC, as nonverbal emotional cues that have a significantly greater effect on moral thinking. Besides, self-identified female participants may perceive the AI assistant’s influence, when displaying SC rather than the baseline, as nonverbal emotional cues that have a significantly greater effect on moral thinking.
@@
We also find that in AC, the perception of AI assistant’ influence on moral thinking in self-identified males is significantly higher than that in self-identified females ($p<0.05$). However, in SC, the perception reflected in self-identified males is significantly lower than in self-identified females ($p<0.01$).
@@ 5.2. The Interaction Effect of Gender and Emotion on Perception
Our results shows that self-identified male participants may perceive the AI assistant’s influence, when displaying AC rather than SC, as nonverbal emotional cues that have a significantly greater effect on moral thinking. However, regardless of how they perceive the AI assistant, the reversal effect on certainty by the AI assistant is the same for participants of both genders, with the reversal effect of AC being more significant. This is an interesting phenomenon where people’s perception of AI does not align with the actual impact it has on them. The reasons behind this discrepancy in perception are worth exploring.
There is a stereotype that different genders tend to express different emotions. Studies have found that emotions such as sadness and depression are often seen as associated with women, while anger and aggression are generally considered to be more closely associated with men (Simon and Nath, 2004; Barrett and Bliss-Moreau, 2009; Craig and Lee, 2020; Plante et al., 2000). Building on our research, which found that participants of different genders perceive AI emotions differently, it raises the question of whether this phenomenon is related to the aforementioned stereotype. In other words, do the emotions that men and women are stereotypically expected to express also influence how they perceive emotions expressed by AI agents? This is an area that can be further explored.
@@ 5.4. Limitations and Future Opportunities
Second, our study focused exclusively on the AI assistant designed to challenge users’ decisions without including AI that supports users or provides detailed analyses. This limitation prevents us from assessing how different AI roles influence user behavior and decision-making. Future research could compare challenging and supportive AI across various scenarios to identify the most appropriate approach for each context.
@@ 6. CONCLUSION
This study empirically examines the influence of nonverbal emotional cues of the AI assistant on decision certainty and how humans perceive the influence of AI in moral dilemmas. Our findings demonstrate that agents expressing anger significantly induce reversal shifts in decision certainty. The interaction effects between participant gender and agents’ nonverbal emotional cues notably affect participants’ perceptions of AI’s influence. These results highlight the potential of subtle nonverbal emotional cues to shape human moral decisions, offering insights for designing future human-AI systems. We hope our findings contribute to ongoing discussions in the AI ethics community about the nuanced ways AI can impact decision-making processes.